# Holographic phase transitions via thermally-assisted tunneling

### v1

@@ 2 Review of the holographic setup
where $M_{5}$ is the 5D Planck scale, $\Lambda_{5}$ is the bulk cosmological constant, ${\cal T}_{i}$ are brane tensions satisfying $-\Lambda_{5}={\cal T}_{\rm UV}=-{\cal T}_{\rm IR}=24M_{5}^{3}$ and $g_{i}$ is the induced metric [^4] Note that in (5) we have not explicitly written the Gibbons–Hawking–York boundary term which is also needed to have a well-defined action with boundary terms Agashe:2020lfz .. The action of the Goldberger–Wise scalar field is
@@
where $\epsilon_{2},\epsilon_{3}<0$. The signs of $\epsilon_{2,3}$ are chosen so that the corresponding deformation in the dual gauge theory is a relevant operator that becomes larger at IR scales. We also assume $0.01\lesssim|\epsilon_{2}|,|\epsilon_{3}|\lesssim 0.1$ as is necessary to obtain the hierarchy $\rho_{\rm IR}\ll\rho_{\rm UV}$. The boundary potentials $U_{i}(\chi)$ are chosen as follows
@@
The potential (8) and the boundary conditions determine the bulk profile $\chi(\rho)$. The profile depends on the potential parameters $\epsilon_{2},\epsilon_{3},v_{\rm UV},\alpha_{\rm IR}$ as well as the parameters of the 5D solution $\rho_{\rm IR}$ and $T$ [^5] One can rescale $\rho$ to set the position of the UV brane to be at $\rho_{\rm UV}=1$. Knowing the profile, one can calculate the 4D effective action for the radion field $\varphi$ (with $\langle\varphi\rangle\equiv\rho_{\rm IR}$),
@@
For convenience, we will work with the parameters $|V_{\rm min}|$ and $m_{\varphi}^{2}$ in what follows, since $|V_{\rm min}|$ is directly related to the critical temperature, as will be explained in section 3, while $m_{\varphi}^{2}$ controls the curvature of the potential and so determines the quantum-to-classical transition temperature $T_{q}$. In particular, we will consider two benchmark values of the radion mass squared: $m_{\varphi}^{2}=0.1$ and $0.9$. These are shown in the right panel of figure 1 and correspond to the parameter values $(\kappa^{4},\epsilon_{2},\epsilon_{3})=(0.25,-0.1,0)$ and $(0.1,-0.04,-0.12)$ of the potential (11), respectively. It is worth noting that the radion mass squared is given by $m_{\varphi}^{2}=-4\kappa^{4}\epsilon_{2}$ for $\epsilon_{3}=0$ and is further enhanced when the cubic coupling $\epsilon_{3}$ is nonzero [^6] It can be numerically checked that $m_{\varphi}^{2}$ also depends on $\alpha_{\rm IR}$ and $v_{\rm UV}$ for non-zero $\epsilon_{3}$.. The remaining parameters, $\alpha_{\rm IR}$, $v_{\rm UV}$, will be adjusted to the desired value of $|V_{\rm min}|$ keeping $m_{\varphi}^{2}$ fixed to the benchmark values. Finally, we assume $N_{c}=3,6$ as the benchmark numbers of colors entering the overall factor in the action (10). [^7] For all numerical computations, $N_{c}$ is first conveniently set to unity and then the obtained action is scaled by the appropriate value of $N_{c}^{2}$.
@@ 3 Euclidean solutions and decay rate
Under this assumption, the problem reduces to finding the solution of the Euclidean equation of motion for $\varphi$ following from the flat-space action (10).999 We neglect the spacetime curvature induced by the false vacuum energy density. This is justified as long as the size of the bounce $\ell_{b}$ is much smaller than the Hubble size in the false vacuum $H\sim T_{c}^{2}/(M_{P}N_{c})$. Below we will see that $\ell_{b}\sim T_{n}^{-1}$, resulting in the bound $T_{n}\gg T_{c}^{2}/(M_{P}N_{c})$ which is always satisfied in our calculations. Assuming spherical symmetry with radial coordinate $r=\sqrt{x_{1}^{2}+x_{2}^{2}+x_{3}^{2}}$, the Euclidean equation of motion for the bounce takes the form
@@
We numerically solve the equation of motion (19) with the boundary conditions (20) and (21). First, let us discuss the time-independent configuration that exists at any temperature $T<T_{c}$. The top row of figure 2 shows this configuration for certain parameters of the radion potential with $m_{\varphi}^{2}=0.9$ and several values of $T$. The solution agrees with the one studied in Ref. Mishra:2023kiu for similar model parameters. Note that away from $T_{c}$ the critical bubble cannot be described using the analytic thin-wall approximation and the numerical solution of eq. 19 is needed.
@@
Before closing this section, let us make an important comment. The temperature $T_{q}$, below which decays via thermal tunneling dominate those via classical transitions, has been identified with the temperature at which the critical bubble develops an extra negative mode. This implies that the thermal bounce describing the tunneling is deformed continuously and merges with the critical bubble in the limit $T\to T_{q}$, and figure 2 seems to suggest that this is the case. However, a closer examination of the bounce solution near $T_{q}$ reveals a more complicated behavior, as discussed in appendix C. It turns out that the thermal bounces do not branch off from the critical bubble at $T=T_{q}$,101010 This behavior is suspected from figure 4 by noticing that the action $B(T)$ is not a smooth function of $T$ at $T=T_{q}$. nor does $T_{q}$ exactly equal the temperature at which the second negative mode appears. The difference between the two temperatures is negligible in practice, and amounts to less than one percent for the case $m_{\varphi}^{2}=0.9$. Nevertheless, it reveals that there are subtleties associated with the holographic bounce that do not typically appear in the usual 4D vacuum decay.
@@ 4 Nucleation temperature
Figure 5: The nucleation temperature $T_{n}$ (scaled by 1/$T_{c}$) as a function of the critical temperature $T_{c}$ for $N_{c}=3$ ($6$) shown as solid blue (red) lines, assuming $m_{\varphi}^{2}=0.1$ (left plot) and $0.9$ (right plot). For comparison, the nucleation temperature determined from the critical bubble solution is shown as a dashed blue (red) line for $N_{c}=3$ ($6$). The horizontal gray dot-dashed line indicates the classical-to-quantum transition temperature, $T_{q}$ (again scaled by $1/T_{c}$) which separates the classical and quantum regimes. The values of the ratio where $T_{n}/T_{c}>1$ are not physically relevant. Note that for each value of $N_{c}$, the curves end at a maximum value of $T_{c}$, above which the transition never takes place $(i.e.~{}\Gamma<H^{4})$.
We find that for small radion mass squared, $m_{\varphi}^{2}=0.1$, corresponding to the absence of nonlinearity in the Goldberger–Wise potential (i.e. $\epsilon_{3}=0$), the above conditions cannot be satisfied for a TeV-scale transition with $N_{c}=3$, and the critical bubble is the relevant solution. This is illustrated on the left panel of figure 5 and agrees with the analysis of Ref. Agashe:2020lfz . On the contrary, for $N_{c}=6$, the thermal bounce solution crucially enhances the vacuum decay rate to complete the phase transition at the TeV scale, which would not be possible using only the critical bubble solution. Moreover, the thermal bounce remains relevant at higher critical temperatures, $T_{c}\gtrsim 10^{6}$ GeV for $N_{c}=3$. Similarly, for large radion mass, $m_{\varphi}^{2}=0.9$, the required conditions can be satisfied already for the TeV scale transition with $N_{c}=6$, as is shown on the right panel of figure 5. It is then important to use the thermal bounce action (22) to find the nucleation temperature correctly. For example, using the critical bubble instead of the thermal bounce would lead to underestimating $T_{n}$ by one order of magnitude at $T_{c}=10^{10}$ GeV and $N_{c}=6$. Note that the endpoints of the functions $T_{n}(T_{c})$ shown in figure 5 correspond to the maximum of the decay rate, beyond which $\Gamma(T)$ is always smaller than $H^{4}$ and no phase transition occurs.
@@ 5 Gravitational wave signal
In this section, we calculate the GW signal that arises from the phase transition mediated by the quantum thermal bounce. We assume that the dominant source of the gravitational waves is due to the collision of bubble walls. This is reasonable, since at large supercooling the ambient radiation is dilute and exerts little pressure on the bubbles which, therefore, can expand in the run-away mode Caprini:2015zlo ; Caprini:2019egz . Upon completion of the phase transition, most of the energy contained in the bubble walls is transferred to the thermal plasma, which attains a temperature $T_{*}\approx T_{c}$.
Figure 6: The abundance of the stochastic GWs as a function of frequency due to collisions of bubbles nucleated via thermal tunneling. The red (blue) lines correspond to $m_{\varphi}^{2}=0.1$ ($0.9$), assuming $g_{*}=100$ and $T_{c}=T_{*}=10^{3}$ GeV (solid lines) and $10^{8}$ GeV (dot-dashed lines). The green, magenta and purple dashed lines show the projected discovery reach from LISA LISA:2017pwj ; Baker:2019nia , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi ; Yagi:2011wg and DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb , respectively.
@@
The GW signal produced by phase transitions dominated by thermally-assisted tunneling is shown in figure 6. As the figure depicts, thermal tunneling does induce a potentially observable signal at future gravitational wave observatories both for the small and large radion mass cases. The small radion mass case ($m_{\varphi}^{2}=0.1$) requires a larger (smaller) number of colors $N_{c}$ at lower (higher) values of $T_{c}$, whereas the larger radion mass case ($m_{\varphi}^{2}=0.9$) entails larger $N_{c}$ at any $T_{c}$. It is particularly interesting that even for high critical temperatures, $T_{c}=10^{8}$ GeV, the signal is within reach of future GW observatories where supercooled phase transitions dominated by tunneling can be probed. Importantly, using only the critical bubble solution would not lead to any GW signal for the TeV-scale phase transition shown in figure 6 with $N_{c}=6$ and $m_{\varphi}^{2}=0.1$. Finally, note that while we only showed signals for the benchmark values in figure 6, the thermal bounce would also give new signals for other parameter values, thereby extending the parameter space in holographic models beyond the range studied with the critical bubble.
@@ 6 Conclusion
Importantly, our solution is calculated in the region of validity of the 4D theory. The full 5D solution – a gravitational instanton interpolating between the two geometries – can, in principle, be constructed, but this is challenging even in the $O(3)$-symmetric case Mishra:2023kiu ; Mishra:2024ehr . Besides, the relevance of such an instanton for vacuum decay is hard to establish, since the negative mode counting argument cannot be literally applied, see Bramberger:2019mkv and references therein. Nevertheless, it is an interesting direction of future research.
@@ Appendix A Negative modes of the critical bubble
$-\bar{\chi}_{n,k}^{\prime\prime}-\frac{2}{\bar{r}}\,\bar{\chi}_{n,k}^{\prime}+\left[\bar{\omega}_{n}^{2}+\frac{\mathrm{d}^{2}\bar{V}(\bar{\varphi}_{0})}{\mathrm{d}\bar{\varphi}^{2}}\right]\bar{\chi}_{n,k}=\lambda^{(k)}_{n}\bar{\chi}_{n,k}\,,$ (33)
@@
To verify this intuition, we solve numerically eq. 35 at different temperatures. At any $T$ we find the ground state $\bar{\chi}_{0,0}(\bar{r})$ and the corresponding eigenvalue $\lambda_{0}^{(0)}$. As shown in figure 7, the function $\bar{\chi}_{0,0}(\bar{r})$ becomes exponentially small at $\bar{r}\approx\bar{r}_{0}$, and therefore the boundary condition (36) does not modify the spectrum. Furthermore, we find that the value $\lambda_{0}^{(0)}$ decreases as the temperature decreases. As soon as it drops below $-4\pi^{2}$, the second negative eigenvalue appears in the dipole sector $n=1$:
@@ Appendix B Numerical calculation of the thermal bounce
In the coordinates ($\tilde{\tau},\tilde{r}$), the integration domain now becomes $10\leqslant\tilde{\tau}<$ 1 and $0\leqslant\tilde{r}\leqslant 1$. The $\tilde{\tau}$ range corresponds to the period $1/T$ in the imaginary time $\tau$, and the $\tilde{r}$ range follows from the boundary location (38). The equation of motion (30) and the boundary conditions (31) in the coordinates (39) are given by
@@
Using the rectangular grid in the $(\tilde{\tau},\tilde{r})$-coordinates, we solve eq. 40 with the boundary conditions (41)-(42) and the condition of periodicity in $\tilde{\tau}$ by the Newton–Raphson method Press:2007ipz to obtain the thermal bounce $\bar{\varphi}_{b}(\tilde{\tau},\tilde{r})$ and its boundary determined by the function $\zeta(\tilde{\tau})$. As an initial guess for the Newton–Raphson method, at temperatures $T\lesssim T_{q}$ we use the critical bubble, which can be easily found by solving the corresponding differential equation, which is perturbed around its extra negative mode (see appendix A). We then obtain the solution iteratively by lowering the temperature and using the bounce solution found at the slightly higher temperature as the new initial guess. The solution evaluated at a particular temperature and an example of the grid used to obtain it are shown in figure 8. Finally, using the variables defined in (29) and the coordinates $(\tilde{\tau},\tilde{r})$, the bounce action (22) becomes
@@
Note that at sufficiently low temperatures the thermal bounce becomes $O(4)$-symmetric on the time interval $10\leqslant\tilde{\tau}<$ 1 and the coordinates $(\tilde{\tau},\tilde{r})$ become inappropriate. This is depicted in the right panel of figure 8 which shows that as the temperature decreases, the boundary changes from being located at a constant $\bar{r}$ value, corresponding to the critical bubble, to being located on the semi-circle $\bar{\tau}^{2}+\bar{r}^{2}=\,$constant, corresponding to the $O(4)$-symmetric solution. Thus, for sufficiently low temperatures it is more appropriate to switch to the radial coordinate $\bar{R}=\sqrt{\bar{\tau}^{2}+\bar{r}^{2}}$. The equation of motion (30) then reduces to an ordinary differential equation, and the boundary value problem (30), (31) is straightforward to solve. The bounce action reads

### v2

@@ 2 Review of the holographic setup
where $M_{5}$ is the 5D Planck scale, $\Lambda_{5}$ is the bulk cosmological constant, ${\cal T}_{i}$ are brane tensions satisfying $\Lambda_{5}={\cal T}_{\rm UV}=-{\cal T}_{\rm IR}=24M_{5}^{3}$ and $g_{i}$ is the induced metric [^4] Note that in (5) we have not explicitly written the Gibbons–Hawking–York boundary term which is also needed to have a well-defined action with boundary terms Agashe:2020lfz .. The action of the Goldberger–Wise scalar field is
@@
where $\epsilon_{2},\epsilon_{3}<0$. The signs of $\epsilon_{2,3}$ are chosen so that the corresponding deformation in the dual gauge theory is a relevant operator that becomes larger at IR scales. We also assume $0.01\lesssim|\epsilon_{2}|,|\epsilon_{3}|\lesssim 0.1$ ensuring that the Goldberger-Wise profile $\chi\left(\rho\right)$ varies slowly across the bulk and remains sensitive to both branes, which is necessary for generating a radion potential capable of stabilizing a large hierarchy: $\rho_{\rm IR}\ll\rho_{\rm UV}$ Goldberger:1999uk ; Mishra:2023kiu . The boundary potentials $U_{i}(\chi)$ are chosen as follows
@@
The potential (8) and the boundary conditions determine the bulk profile $\chi(\rho)$. The profile depends on the potential parameters $\epsilon_{2},\epsilon_{3},v_{\rm UV},\alpha_{\rm IR}$ as well as the parameters of the 5D solution $\rho_{\rm IR}$ and TT. [^5] One can rescale $\rho$ to set the position of the UV brane to be at $\rho_{\rm UV}=1$. Knowing the profile, one can calculate the 4D effective action for the radion field $\varphi$ (with $\langle\varphi\rangle\equiv\rho_{\rm IR}$),
@@
For convenience, we will work with the parameters $|V_{\rm min}|$ and $m_{\varphi}^{2}$ in what follows, since $|V_{\rm min}|$ is directly related to the critical temperature, as will be explained in section 3, while $m_{\varphi}^{2}$ controls the curvature of the potential and so determines the quantum-to-classical transition temperature $T_{q}$. In particular, we will consider two benchmark values of the radion mass squared: $m_{\varphi}^{2}=0.1$ and $0.9$. These are shown in the right panel of figure 1 and correspond to the parameter values $(\kappa^{4},\epsilon_{2},\epsilon_{3})=(0.25,-0.1,0)$ and (0.1,0.04,0.12)(0.1,-0.04,-0.12) of the potential (11), respectively. It is worth noting that the radion mass squared is given by $m_{\varphi}^{2}=-4\kappa^{4}\epsilon_{2}$ for $\epsilon_{3}=0$ and is further enhanced when the cubic coupling $\epsilon_{3}$ is nonzero [^6] It can be numerically checked that $m_{\varphi}^{2}$ also depends on $\alpha_{\rm IR}$ and $v_{\rm UV}$ for non-zero $\epsilon_{3}$.. The remaining parameters, $\alpha_{\rm IR}$, $v_{\rm UV}$, will be adjusted to the desired value of $|V_{\rm min}|$ keeping $m_{\varphi}^{2}$ fixed to the benchmark values. Finally, we assume $N_{c}=3,6$ as the benchmark numbers of colors entering the overall factor in the action (10). [^7] For all numerical computations, $N_{c}$ is first conveniently set to unity and then the obtained action is scaled by the appropriate value of $N_{c}^{2}$.
@@ 3 Euclidean solutions and decay rate
Under this assumption, the problem reduces to finding the solution of the Euclidean equation of motion for $\varphi$ following from the flat-space action (10). [^9] We neglect the spacetime curvature induced by the false vacuum energy density. This is justified as long as the size of the bounce $\ell_{b}$ is much smaller than the Hubble size in the false vacuum $H\sim T_{c}^{2}/(M_{P}N_{c})$. Below we will see that $\ell_{b}\sim T_{n}^{-1}$, resulting in the bound $T_{n}\gg T_{c}^{2}/(M_{P}N_{c})$ which is always satisfied in our calculations. Assuming spherical symmetry with radial coordinate $r=\sqrt{x_{1}^{2}+x_{2}^{2}+x_{3}^{2}}$, the Euclidean equation of motion for the bounce takes the form
@@
We numerically solve the equation of motion (19) with the boundary conditions (20) and (21). First, let us discuss the time-independent configuration that exists at any temperature $T<T_{c}$. The top row of figure 2 shows this configuration for certain parameters of the radion potential with $m_{\varphi}^{2}=0.9$ and several values of $T$. The solution agrees with the one studied in Ref. Mishra:2023kiu for similar model parameters. Note that away from $T_{c}$ the critical bubble cannot be described using the analytic thin-wall approximation and the numerical solution of eq. 19 is needed. [^10] At sufficiently low temperatures $T\ll T_{c}$, when the true vacuum lies significantly deeper than the potential barrier, the thick-wall approximation may be useful. However, we prefer to use the full numerical bounce solution which is applicable across the full temperature range considered.
@@
Before closing this section, let us make an important comment. The temperature $T_{q}$, below which decays via thermal tunneling dominate those via classical transitions, has been identified with the temperature at which the critical bubble develops an extra negative mode. This implies that the thermal bounce describing the tunneling is deformed continuously and merges with the critical bubble in the limit $T\to T_{q}$, and figure 2 seems to suggest that this is the case. However, a closer examination of the bounce solution near $T_{q}$ reveals a more complicated behavior, as discussed in appendix C. It turns out that the thermal bounces do not branch off from the critical bubble at $T=T_{q}$, [^11] This behavior is suspected from figure 4 by noticing that the action $B(T)$ is not a smooth function of $T$ at $T=T_{q}$. nor does $T_{q}$ exactly equal the temperature at which the second negative mode appears. The difference between the two temperatures is negligible in practice, and amounts to less than one percent for the case $m_{\varphi}^{2}=0.9$. Nevertheless, it reveals that there are subtleties associated with the holographic bounce that do not typically appear in the usual 4D vacuum decay.
@@ 4 Nucleation temperature
[…]
We find that for small radion mass squared, $m_{\varphi}^{2}=0.1$, corresponding to the absence of nonlinearity in the Goldberger–Wise potential (i.e. $\epsilon_{3}=0$), the above conditions cannot be satisfied for a TeV-scale transition with $N_{c}=3$, and the critical bubble is the relevant solution. This is illustrated on the left panel of figure 5 and agrees with the analysis of Ref. Agashe:2020lfz . On the contrary, for $N_{c}=6$, the thermal bounce solution crucially enhances the vacuum decay rate to complete the phase transition at the TeV scale, which would not be possible using only the critical bubble solution. Moreover, the thermal bounce remains relevant at higher critical temperatures, $T_{c}\gtrsim 10^{6}$ GeV for $N_{c}=3$. Similarly, for large radion mass, $m_{\varphi}^{2}=0.9$, the required conditions can be satisfied already for the TeV scale transition with $N_{c}=6$, as is shown on the right panel of figure 5. It is then important to use the thermal bounce action (22) to find the nucleation temperature correctly. For example, using the critical bubble instead of the thermal bounce would lead to underestimating $T_{n}$ by one order of magnitude at $T_{c}=10^{10}$ GeV and $N_{c}=6$. To get the correct value of $T_{n}$ using the critical bubble, one would have to shift the radion mass by a factor of $\mathcal{O}(1)$. Note that the endpoints of the functions $T_{n}(T_{c})$ shown in figure 5 correspond to the maximum of the decay rate, beyond which $\Gamma(T)$ is always smaller than $H^{4}$ and no phase transition occurs.
Figure 6: The ratio $\beta/H$, defined in (28), as a function of the critical temperature $T_{c}$ for $N_{c}=3$ (66) shown as solid blue (red) lines, assuming $m_{\varphi}^{2}=0.1$ (left plot) and $0.9$ (right plot). For comparison, the value of $\beta/H$ determined from the subdominant critical bubble solution is shown as a dashed blue (red) line for $N_{c}=3$ (66). In the left plot, the solid blue line exhibits a discontinuity near $T_{c}=10^{6}$ GeV, which corresponds to where the thermal bounce branches off from the critical bubble (see figure 4).
@@ 5 Gravitational wave signal
In this section, we calculate the GW signal that arises from the phase transition mediated by the quantum thermal bounce. We assume that the dominant source of the gravitational waves is due to the collision of bubble walls. This is reasonable, since at large supercooling the ambient radiation is dilute and exerts little pressure on the bubbles which, therefore, can expand in the run-away mode Caprini:2015zlo ; Caprini:2019egz . Upon completion of the phase transition, most of the energy contained in the bubble walls is transferred to the thermal plasma, which attains a temperature $T_{*}\approx T_{c}$ ; this is a good approximation for a strong first-order transition where the large latent heat reheats the universe.
Figure 7: The fractional energy density of the stochastic GWs as a function of frequency due to collisions of bubbles nucleated via thermal tunneling. The red (blue) lines correspond to $m_{\varphi}^{2}=0.1$ ($0.9$), assuming $g_{*}=100$ and $T_{c}=T_{*}=10^{3}$ GeV (solid lines) and $10^{8}$ GeV (dot-dashed lines). The blue dotted line shows the GW signal as predicted by the static critical bubble solution, which compares with the true time-dependent bounce solution (blue solid line) for the same parameters. The green, magenta and purple dashed lines show the projected discovery reach from LISA LISA:2017pwj ; Baker:2019nia , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi ; Yagi:2011wg and DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb , respectively.
@@
In Fig. 6, we show $\beta/H$ as a function of the critical temperature $T_{c}$ obtained using our quantum thermal bounce, and for $T_{n}<T_{q}$, the $\beta/H$ values are compared with those obtained from the (inaccurate) critical bubble solution.
@@
The GW signal produced by phase transitions dominated by thermally-assisted tunneling is shown in figure 7. As the figure depicts, thermal tunneling does induce a potentially observable signal at future gravitational wave observatories both for the small and large radion mass cases. The small radion mass case ($m_{\varphi}^{2}=0.1$) requires a larger (smaller) number of colors $N_{c}$ at lower (higher) values of $T_{c}$, whereas the larger radion mass case ($m_{\varphi}^{2}=0.9$) entails larger $N_{c}$ at any $T_{c}$. It is particularly interesting that even for high critical temperatures, $T_{c}=10^{8}$ GeV, the signal is within reach of future GW observatories where supercooled phase transitions dominated by tunneling can be probed. Importantly, using only the critical bubble solution would not lead to any GW signal for the TeV-scale phase transition shown in figure 7 with $N_{c}=6$ and $m_{\varphi}^{2}=0.1$. Similarly, for larger radion mass $m_{\varphi}^{2}=0.9$, again with $N_{c}=6$, the GW signal predicted by the thermal tunneling solution (blue solid line) differs by an order of magnitude from that of the critical bubble solution (blue dotted line) [^12] However, note that readjusting the radion mass by an $\mathcal{O}(1)$ factor can reproduce the same signal obtained from the (inaccurate) critical bubble solution.. Finally, note that while we only showed signals for the benchmark values in figure 7, the thermal bounce would also give new signals for other parameter values, thereby extending the parameter space in holographic models beyond the range studied with the critical bubble.
@@ 6 Conclusion
Importantly, our solution is calculated in the region of validity of the 4D theory. The full 5D solution – a gravitational instanton interpolating between the two geometries – can, in principle, be constructed, but this is challenging even in the $O(3)$-symmetric case Mishra:2023kiu ; Mishra:2024ehr . Besides, the relevance of such an instanton for vacuum decay is hard to establish, since the negative mode counting argument cannot be literally applied, see Bramberger:2019mkv and references therein. Nevertheless, it is an important direction of future research, since it will enable a more accurate determination of the bounce action.
@@ Appendix A Negative modes of the critical bubble
$\bar{\chi}_{n,k}^{\prime\prime}-\frac{2}{\bar{r}}\,\bar{\chi}_{n,k}^{\prime}+\left[\bar{\omega}_{n}^{2}+\frac{\mathrm{d}^{2}\bar{V}(\bar{\varphi}_{0})}{\mathrm{d}\bar{\varphi}^{2}}\right]\bar{\chi}_{n,k}=\lambda^{(k)}_{n}\bar{\chi}_{n,k}\,$, (33)
@@
To verify this intuition, we solve numerically eq. 35 at different temperatures. At any $T$ we find the ground state $\bar{\chi}_{0,0}(\bar{r})$ and the corresponding eigenvalue $\lambda_{0}^{(0)}$. As shown in figure 8, the function $\bar{\chi}_{0,0}(\bar{r})$ becomes exponentially small at $\bar{r}\approx\bar{r}_{0}$, and therefore the boundary condition (36) does not modify the spectrum. Furthermore, we find that the value $\lambda_{0}^{(0)}$ decreases as the temperature decreases. As soon as it drops below $4\pi^{2}$, the second negative eigenvalue appears in the dipole sector $n=1$:
@@ Appendix B Numerical calculation of the thermal bounce
In the coordinates ($\tilde{\tau},\tilde{r}$), the integration domain now becomes $0\leqslant\tilde{\tau}<1$ and $0\leqslant\tilde{r}\leqslant 1$. The $\tilde{\tau}$ range corresponds to the period $1/T$ in the imaginary time $\tau$, and the $\tilde{r}$ range follows from the boundary location (38). The equation of motion (30) and the boundary conditions (31) in the coordinates (39) are given by
@@
[…]
@@
[…]